Calculate the pH and degree of dissociation of:-
a. 0.8M HF
b. The same solution which is also in 0.1M NaF, If ka=6.8x10-4.
1
Expert's answer
2020-09-21T06:31:54-0400

HF --> H+ + F-


a) Ka ≈ [H+][F-]/[HF]

[H+] = [F-]

[H+]2 = Ka[HF] = 6.8*10-4 * 0.8 = 5.44*10-4

[H+] = 2.3*10-2 M

pH = -log [H+] = 1.64


"K_a=\alpha^2C/(1-\alpha)"

6.8*10-4 = "(\alpha^2*0.8)/(1-\alpha)"

"\alpha=0.029" (or 2.9 %)


b) For buffer solutions

pH = pKa + log [salt]/[acid]

Therefore,

pH = pKa + log [NaF]/[HF]

pKa = -log Ka

pH = -log Ka + log [NaF]/[HF] = 3.17 + log (0.1 / 0.8) = 3.17 - 0.90 = 2.27


Answer:

"\alpha=0.029" (or 2.9 %)

a) pH = 1.64

b) pH = 2.27